Vocaloid:

Find the line integral of f(x,y) = ye^(x^2) along the curve r(t) = -4t i - 3t j from t = -2 to t = 0

Vocaloid:

Please do not comment unless you are actually familiar with the subject and are making a genuine attempt to help.

sillybilly123:

\(\int ~ ds ~ ~ ~ \{y e ^ {x^2} \}\) By Def: \( ds = \sqrt{dx^2 + dy ^2 } ~ \cdot ~ dt\) \(x= - 4 t, ~ dx = - 4 ~ dt\) \(y= - 3 t, ~ dy = - 3 ~ dt\) \( ds = \sqrt{16 + 9 } ~ dt = 5 ~ dt\) \(\implies 5 \int ~ dt ~ ~ (- 3t) e ^ {16 t^2} \) \(= - \dfrac{15}{32}\left[ e ^ {16 t^2} \right]_{-2}^{0}\) \(= \dfrac{15}{32} e ^ {64} \) if that's rubbish, do holler :)

Vocaloid:

thank you C:

sillybilly123:

Ouch. By Def 🤨🥴 \(ds = \sqrt{\dot x^2 + \dot y ^2 } ~ \cdot ~ dt ~ ~ \because ds^2 = dx^2 + dy^2\)